OpenStudy (anonymous):

If tan x° = 3 divided by y and cos x° = y divided by z ,what is the value of sin x°?

OpenStudy (bahrom7893):

tan(x) = 3/y cos(x) = y/z tan(x) = sin(x)/cos(x) 3/y = sin(x)/(y/z)

OpenStudy (bahrom7893):

multiply both sides by y/z: (y/z)*(3/y) = sin(x) Sin(x) = 3/z